OpenStudy (anonymous):

Find dy/dx in terms of x and y if x^(1/3)=2sqrt(y)

OpenStudy (anonymous):

\[x \sqrt[3]{x}=2\sqrt{y}\]

OpenStudy (anonymous):

you can just solve this for y and directly differentiate... no real need for implicit differentiation

OpenStudy (radar):

\[\sqrt[3]{x ^{4}}=2\sqrt{y}\]

OpenStudy (radar):

\[x ^{4/3}=2y ^{1/2}\]

OpenStudy (radar):

\[y ^{1/2}=.5x ^{4/3}\]

OpenStudy (radar):

\[y=(.5x ^{4/3})^{1/2}\]

OpenStudy (radar):

\[y=.5^{1/2}x ^{2/3}\]

OpenStudy (anonymous):

since the ?says ito of y and x so implicit is important \[x^{1/3}=2\sqrt{y}\] \[\frac{ 1 }{ 3x^{2/3} }=\frac{- y' }{ \sqrt{y} }\] \[y'=-\frac{ \sqrt{y} }{ 3\sqrt[3]{x^2} }\] in terms of x and y